OpenStudy (anonymous):

-0.5(3y-4)+0.3y=0.2(15-y)

OpenStudy (anonymous):

-1.5y + 2 + 0.3y = 3 - 0.2y (-1.5+0.3+0.2) y = 1 -1y = 1 y = -1

OpenStudy (anonymous):

-1.5y + 2 + 0.3y = 3 - 0.2y -y = 1 y=-1

OpenStudy (anonymous):

-0.5(3y-4)+0.3y=0.2(15-y) ===> - 1.5y+2+0.3y =3-0.2y ===> - y =1 ==> y =-1